Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
berni - 15:52 Wednesday 18 February 2026 (68709) Print this report
Operator Report - Morning shift

ITF found unlocked: west cavity locked and north cavity unlocked. To recover the lock I had to move the PR top stage. After that the ITF was locked in LN2 and then in LN3_Aligned at 8:13 UTC.

Once locked Romain and Alba started working at the OMC mode matching.

At 11:23 UTC the ITF was manually unlocked to continue the activity in NI single bounce with PR TY further misaligned of -30ur.

Activity still in progress at the end of the shift.

Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×